A balloon starting from the ground has been ascending vertically at a uniform velocity for 4 seconds and a stone let fall from it reaches the ground in 6 seconds . find the velocity of the balloon and its height when the stone was let fall (g=10m/sec sq)

Grade:11

1 Answers

Shreya dadhich
5 years ago
the block is moving with uniform speed 
t=4s, a=-10 m/s^2, u=0 m/s 
v=u+at
v=-40 m/s 
The speed of balloon is 40m/s upwards 
for the stone
u=-40m/s, t=6s ,∆y=ut+1/2 at^2
∆y=-60m 
The height is 60 m up from the ground